Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-02-14
ISS On-Orbit Status 02/14/08

Mission 1E Flight Day 8 (FD8).

Crew sleep cycle shifted one hour earlier: 3:45am - 7:15pm for both crews.

Last night, Mission 1E was extended by one day (13+0+2), with landing now on Wednesday, 2/20, at ~9:03am EST (if at KSC).

Columbus final activation has been completed.

After yesterday's command queue lockup between the COL CCS (Columbus Orbital Laboratory Command & Control System) and the COL MMC (Mission Management Computer), final activation of COL data management systems was accomplished late last night by transitioning/swapping the primary & backup C&C MDM (Multiplexer/Demultiplexer) computers. (The MMC is the intermodule interface computer required between the US C&DH (Command & Data Handling) system and the COL DMC (Data Management Computer) which handles equipment monitoring)

With IMV (Intermodular Ventilation) and smoke detector function established, the crew continued COL outfitting and activation, including:

Activation of the PWS (Portable Workstation) computer,
Installation & activation of PCS (Portable Computer System) laptops,
Relocation of the Biolab (BLB) from its launch location to its ISPR (International Standard Payload Rack) plus installing its DC converter, and
Installation of VCA (Video Camera Assembly) 1 & 2.
Other outfitting activities completed by last night ahead of timeline included -

Unlocking the anti-vibration mounts for the WPA (Water Pump Assembly) and CWSA (Condensate Water Separator Assembly),
Relocating the EDR (European Drawer Rack) from its launch position to its operational location,
Removing anti-vibration bolts on the CFA (Cabin Fan Assembly),
Installing a ZSR (Zero-G Stowage Rack) in COL, and
Rotating the COL D1 Rack to gain access to payload stowage locations and recording a video of the activities for subsequent downlink to COL-CC (Columbus Orbital Laboratory Control Center) in Oberpfaffenhofen).
(Biolab (Biological Experiment Laboratory in Columbus) is a biological and physiological facility for studying the effects of microgravity on microorganisms, cells, tissue cultures, small plants and small invertebrates. EDR is a multi-discipline facility that provides an opportunity for compact experiments of low mass, supporting a variety of small payloads. Each payload will have access to power, data communications, cooling among other accommodations.)

Today, CDR Whitson completed COL vestibule outfitting by installing the nitrogen (N2) supply jumper/hose between Node-2 and COL, followed by purge of the N2 system to remove any contamination from the lines.

To prepare for providing COL power to the externally mounted MISSE (Materials ISS Experiment) payload coming up on Mission 1J/A, Leo Eyharts swapped a 20 amp fuse with a 3 amp fuse in the COL PPSB (Payload Power Switch Box).

Before breakfast, for the last time FE-2 Dan Tani completed his daily access of the SLEEP experiment (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) software for data logging and filling in questionnaire entries in the experiment's laptop session file on the HRF-1 laptop for downlink. (To monitor the crewmember's sleep/wake patterns and light exposure, Dan wears a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

For today's recording and ground-commanded downlinking of onboard video, Peggy Whitson inserted and swapped tapes in the two VCRs (Video Cassette Recorders), 1&2.

Whitson had ~1hr set aside for troubleshooting the failed video connection for the ORB-1 video channel from the Shuttle to ISS via the ISS CVIU-10 (Common Video Interface Unit #10) in Node-2. (For the IFM (In-flight Maintenance), the CDR accessed the base of Harmony's S4 rack and used pin kit jumpers to cross-wire CVIU-10 video transmit/receive cables for continuity testing. If this test was unsuccessful, Peggy was to extend the testing to other parts of the cabling.)

Later, Peggy performed a pre-EVA video check by routing a Shuttle view to a Lab RWS (Robotic Workstation) monitor and checking picture quality.
FE-1 Malenchenko hooked up the necessary 'plumbing' and compressor for transferring potable water from the BV2 tank of Progress M-63/28P to the SM Rodnik BV2 water tank. (The water transfer, monitored by Yuri several times during the day, at first was by self-flow (under its own tank pressure), then using the compressor pump via a GZhS gas/liquid separator, to remove air bubbles in the water. Filling of the empty Progress BV1 & BV2 tanks with urine will be scheduled later.)

Starting a new round of RS ventilation system maintenance, the FE-1 worked in the Service Module (SM), replacing the four dust collector filters (PF1-4) and cleaning 'Group A' fan grilles.

For tomorrow's EVA-3 by MS2/EV1 Walheim and MS4/EV3 Love, final preparations by Tani, Rex & Stan today included -

Reconfiguring two DCS 760 digital still cameras for the spacewalk, leaving them connected to station power for now to conserve batteries, one with 28mm lens and a new EVA Action Finder unit, the other with 35mm lens and an EVA flash,
Setting up three batteries in the A/L BSA (Airlock Battery Stowage Assembly) for charging during the prebreathe period and to be installed in the photo equipment prior to egress,
Starting EMU battery charge,
Powering down onboard amateur (ham) radio equipment (Kenwood in SM, Ericsson in FGB) to prevent RF (radio frequency) interference with the EMU (Extravehicular Mobility Unit)/spacesuits,
Preparing the A/L EL (Equipment Lock) for the Campout & spacewalk,
Closing out METOX regeneration in the A/L bake-out oven, and
Reviewing EVA-3 procedures particulars.
Peggy also prepared a special EVA swatch tool for a late-added touch inspection of an A/L handrail (#506) with a possibly sharp edged MMOD impact spot. (The touch tests, supported by close-up photography, will use four swatch samples stroked firmly over the area of interest, once in each of four directions - RTV (Room Temperature Vulcanizing) material and Vectran material on the swatch tool plus RTV and Vectran on a surplus overglove (on hand).)

Whitson, Tani, Malenchenko and the two spacewalkers underwent the standard PMC (Private Medical Conference) via S- & Ku-band audio/video.

Dan Tani closed the Lab science window shutters as protection during today's Atlantis thruster firings that set up attitude for the periodic water dump and also to protect the window during for tomorrow's EVA.

The FE-1 continued the extended leak checking of the spare BZh Liquid Unit (#056) for the Elektron O2 generator by checking the unit's pressure and charging it once again with pressurized N2 from the BPA-M Nitrogen Purge Unit (#23) to 1 atm (1 kg/cm2). The last test pressurization was on 1/09. (During Elektron operation, the inert gas locked up in the BZh has the purpose to prevent dangerous O2/H2 mixing. A leaking BZh cannot be used.)

Yuri also conducted the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS hatch openings (8) in the SM, FGB and DC1 (Docking Compartment).

In preparation for upcoming experiment activities, Malenchenko continued the current extended performance check of the Russian KRIOGEM-03M refrigerator system, powered up on 2/12 to -22 degC with an AX ice pack from storage installed, for once-daily monitoring of its displayed temperature for the next five days (through 2/17).

Yuri also completed the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.)

Later, the FE-1 handled the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

At ~9:55am EST, the crew conducted a 20-min interactive TV 'voice call' event with attendees at an ESA VIP gathering at the German Federal Ministry of Economics & Technology in Berlin, headed by Angela Merkel, the Chancellor of the Federal Republic of Germany, Jean-Jacques Dordain, ESA Director General, and Thomas Reiter as Moderator. The linkup and exchange, focused on the commissioning of the European Columbus laboratory and the resulting increasing salutary 'internationalization' of the ISS, were very successful.

At ~11:10am, the ISS crew also engaged in three live TV/PD-100 camcorder media interviews of about 6.5 min duration each, with NBC News (Tom Costello), WOI-TV (Elizabeth Erwin), and WBBM Radio (Steve Grzanich).

FE-2 Tani and FE-2-16 Eyharts had over 2.5 hrs scheduled top continue their handover activities as ISS 2nd flight engineers.

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR), TVIS treadmill (FE-1, FE-2), RED resistive exercise device (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Malenchenko later transferred the crew's exercise data file to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Rex Walheim (EV1) & Stan Love (EV3) will begin their 'campout' (nachalo desaturatsiy = desaturation start) in the A/L with hatch closure and depressurization of the Crewlock (CL) from 14.7 to 10.2 psi at ~6:00pm (i.e., an hour earlier than yesterday), followed by mask prebreathe at ~6:00-7:15pm. Sleep time for the ISS crew begins at 7:15pm.

Following the usual hygiene break/with mask prebreathe for Rex & Stan at ~4:20-5:30am tomorrow morning after spending the night on 10.2 psi, the A/L hatch will be closed again for EVA preps in 10.2 psi, followed by EMU purge and prebreathe in the EMUs. Afterwards, with CL depressurization and EV1/EV3 egress, EVA-3 nominally begins at ~8:35am EST. Alan Poindexter will support the spacewalk as IV (Intravehicular) crewmember, keeping tabs with the detailed activity steps and crib sheet, and Melvin, Tani & Eyharts will work the SSRMS (Space Station Remote Manipulator System) for the hardware transfers (SOLAR, CMG, EuTEF) . (Objectives of the spacewalk are to transfer SOLAR (Solar Monitoring Observatory) and the EuTEF (European Technology Exposure Facility) to the COL EPF (External Payload Facility), and to retrieve the failed CMG (Control Moment Gyro) for stowage in the Shuttle PLB for return. Get-ahead tasks for EVA-3 include more inspection of the failed SARJ (Solar Alpha Rotary Joint) with close-up photography, inspection of a suspected damage divot or debris (with its removal with a wipe) and photography of the remaining 5 single covers that have not yet been inspected, with 4 of them to be sampled with a sticky tape.)

Transfers: As of this morning, 7 filled CWCs (Contingency Water Containers, contents ~45 L each) and 4 PWRs (Payload Water Reservoirs) were transferred from the Shuttle to ISS. Nitrogen transfer was completed, with 27 lbs N2 transferred to the station. Even with the additional mission extension by one day, 60 lbs of oxygen (O2) will be available for transfer on FD10.

No CEO photo targets uplinked for today.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use